
Chip N10 has the same problem as last week which emerged in the previous night during the second half of the night while DES observations were being taken. The chip could not be repaired for tonight.
4 Dome occolusion warnings went off during the first half of the night but the images were not affected by this. The affected exposures were 771378, 771451, 771462 and 771483.
Second half of night (DES observers): no incidents.
| Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
|---|---|---|---|---|---|---|
| 2018-09-09 | 11.0 | 0.0 | 0.0 | 0.0 | 0.0 | 11 |
Links
[1] mailto:djschelgel@lbl.gov
[2] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/7199/delete?destination=print/node/14490
[3] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/7199/edit?destination=print/node/14490
[4] mailto:frieman@fnal.gov
[5] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/7200/delete?destination=print/node/14490
[6] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/7200/edit?destination=print/node/14490
[7] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/add/node/14490?destination=print/node/14490
file:///usr/remote/user/sispi/nightsum-reports/nightsum-2018-09-09/nightsum.html